# Stag-1: Towards Realistic 4D Driving Simulation with Video Generation Model

### v1

@@ 1 Introduction
Mainstream research increasingly relies on simulation software for extensive algorithm testing and validation [38]. Yet, simulations based on 3D modeling struggle to accurately replicate realistic driving scenarios, creating a substantial gap between synthetic environments and real-world conditions [58]. To address this, current autonomous driving testing solutions strive to build highly realistic scenarios for validating driving algorithms [43, 12]. With the rapid advancement of text-to-image and text-to-video generative models [52], some research has focused on generating trajectory-controlled images or videos to simulate autonomous driving scenes, guided by maps and surrounding vehicle poses to improve scene accuracy [23, 10]. However, real-world driving involves constantly moving pedestrians, vehicles, and objects that introduce structural changes to the environment. Video generation methods often struggle to capture these dynamic changes or the close interactions between elements, leading to inconsistencies in scene continuity, such as background and vehicle type shifts, which complicate maintaining temporal consistency [55, 16]. Recently, approaches based on NeRF [44] and 3D Gaussian Splatting (3DGS) [19, 11] have aimed to capture dynamic elements with greater precision by rendering and modeling 3D scenes. Nonetheless, these methods still face challenges in reconstructing 4D scenes from arbitrary viewpoints, handling extensive dynamic view changes with significant camera movement, and managing long-term temporal transformations under static views.
@@ 3.2 Spatial-Temporal Aligned Reconstruction
D32 PSNR↑ 28.47 28.24 28.16 31.35 32.91 25.14 26.61 25.14 27.44 28.08
SSIM↑ 0.876 0.866 0.806 0.911 0.934 0.813 0.796 0.747 0.857 0.864
LPIPS↓ 0.136 0.252 0.228 0.106 0.104 0.165 0.305 0.313 0.137 0.129
D32 PSNR*↑ 23.26 23.37 24.32 26.02 27.71 20.48 22.21 23.49 22.92 24.98
SSIM*↑ 0.716 0.701 0.682 0.783 0.796 0.753 0.697 0.660 0.680 0.644
S32 PSNR↑ 29.42 28.31 30.00 30.73 31.03 26.82 27.63 28.89 27.05 28.56
SSIM↑ 0.891 0.879 0.834 0.883 0.897 0.836 0.848 0.814 0.825 0.877
LPIPS↓ 0.118 0.196 0.201 0.116 0.112 0.134 0.193 0.212 0.142 0.132
@@ 3.3 Point-Conditioned Video Generation
$\mathrm{Att}\left(\cdot\right)=\begin{cases}\mathrm{S}\-\mathrm{A}(Q_{\mathrm{s}},K_{\mathrm{s}},V_{\mathrm{s}}),&Q_{\mathrm{s}},K_{\mathrm{s}},V_{\mathrm{s}}\in\mathbb{R}^{(B\times T)\times H\times W\times 6}\\ \mathrm{T}\-\mathrm{A}(Q_{\mathrm{t}},K_{\mathrm{t}},V_{\mathrm{t}}),&Q_{\mathrm{t}},K_{\mathrm{t}},V_{\mathrm{t}}\in\mathbb{R}^{(B\times 6)\times H\times W\times T}\end{cases}\vspace{-1mm}$ (13)
Our method captures spatial relationships within each frame across different perspectives, while also considering temporal connections between consecutive time steps. The loss function for this Stage is designed similarly to equation 12.
@@
PSNR↑ 29.64 28.31 31.37 32.34 34.96 34.61 35.23
SSIM↑ 0.918 0.862 0.904 0.886 0.945 0.950 0.954
LPIPS↓ 0.117 0.346 0.246 0.142 0.068 0.05 0.043
PSNR*↑ 16.48 19.55 23.07 25.71 25.46 25.78 26.42
@@ 4.1 Implementation Details
Methods Frozen time Frozen space Remove sars
FID↓ FVD↓ FID↓ FVD↓ FID↓ FVD↓

### v2

@@ 1 Introduction
Mainstream research increasingly relies on simulation software for extensive algorithm testing and validation [38]. Yet, simulations based on 3D modeling struggles to accurately replicate realistic driving scenarios, creating a substantial gap between synthetic environments and real-world conditions [58]. To address this, existing autonomous driving testing solutions strive to build highly realistic scenarios for validating driving algorithms [43, 12]. With the rapid advancement of text-to-image and text-to-video generative models [52], some research has focused on generating trajectory-controlled images or videos to simulate autonomous driving scenes, guided by maps and surrounding vehicle poses to improve scene accuracy [23, 10]. However, real-world driving involves constantly moving pedestrians, vehicles, and objects that introduce structural changes to the environment. Video generation methods often struggle to capture these dynamic changes or the close interactions between elements, leading to inconsistencies in scene continuity, such as background and vehicle type shifts, which complicate maintaining temporal consistency [55, 16]. Recently, approaches based on NeRF [44] and 3D Gaussian Splatting (3DGS) [19, 11] have aimed to capture dynamic elements with greater precision by rendering and modeling 3D scenes. Nonetheless, they still face challenges in reconstructing 4D scenes from arbitrary viewpoints, handling extensive dynamic view changes with significant camera movement, and managing long-term temporal transformations under static views.
@@ 3.2 Spatial-Temporal Aligned Reconstruction
[…]
D32 PSNR*$\uparrow$ 23.26 23.37 24.32 26.02 27.71 20.48 22.21 23.49 22.92 24.98
SSIM*$\uparrow$ 0.716 0.701 0.682 0.783 0.796 0.753 0.697 0.660 0.680 0.644
[…]
@@ 3.3 Point-Conditioned Video Generation
$\mathrm{Att}\left(\cdot\right)=\begin{cases}\mathrm{S}\-\mathrm{A}(Q_{\mathrm{s}},K_{\mathrm{s}},V_{\mathrm{s}}),&Q_{\mathrm{s}},K_{\mathrm{s}},V_{\mathrm{s}}\in\mathbb{R}^{(B\times T)\times H\times W\times 6},\\ \mathrm{T}\-\mathrm{A}(Q_{\mathrm{t}},K_{\mathrm{t}},V_{\mathrm{t}}),&Q_{\mathrm{t}},K_{\mathrm{t}},V_{\mathrm{t}}\in\mathbb{R}^{(B\times 6)\times H\times W\times T}.\end{cases}\vspace{-1mm}$ (13)
The proposed Stag-1 captures the spatial relationships within each frame across different perspectives and also considers temporal correlations between consecutive time steps. The training objective for this stage is designed similarly to 12.
@@
[…]
PSNR*$\uparrow$ 16.48 19.55 23.07 25.71 25.46 25.78 26.42
@@ 4.1 Implementation Details
Methods Frozen time Frozen space Remove cars
[…]